OpenStudy (anonymous):

Fm(x)=x^3-mx+m-2 Prove that all the curves Intersect at a constant point

OpenStudy (anonymous):

you should find derivative. Fm'(x)=3x^2-m then it must equal 0 3x^2-m=0 3x^2=m x^2=m/3 x=sqrt(m/3) or x=-sqrt(m/3) all values of x are constant points. do you understand??

OpenStudy (anonymous):

never understand

OpenStudy (anonymous):

Fm(x)=x^3-mx+m-2 =x^3+m(1-x)-2 let x=1 to get f(1)=3 i think that the point constant is A(1.3)

OpenStudy (anonymous):

i think you have to study :)

OpenStudy (anonymous):

how :) what do you mean

OpenStudy (anonymous):

what is this? lol